The time constant of LR circuit is 20 s.when a resistance of 2Ω is connected in series with R then time constant becomes 2s ,self inductance of circuit is ___

Grade:12

1 Answers

Vikas TU
7 years ago
t = 20 = L/R
and
after adding a resistance of 2 ohm,
t’ = L/(R +  2) = 2
dividing t/t’ we get,
(R+2)/R = 20/2 = 10
R = ¼ ohm and
L = 5 H